Operation Summary for July 1, 2011 - July 4, 2011 |
The Integrated Luminosity for the period from 6/27/11 to 7/4/11 was 59.43 pb-1 (inverse picobarns). NuMI reported receiving 5.12E18 protons on target during this same period. |
Friday July 1 The day shift began Operations monitoring the Tevatron (TEV) store 8854 (54.94E30) and the Recycler stash (197.0E10), with the Antiproton Source (Pbar) stacking (21.0E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-1015 taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8854 at 11:56 AM and put the TEV into shot setup. Operators reported at 12:18 PM that the TEV sector E48 kicker tripped off several times and then would not reset. A TEV kicker person investigated, found nothing wrong, and reset the kicker. Operations established store 8855 at 1:44 PM with an average initial luminosity (AIL) of 304.3E30. Operators turned off the NuMI beam and prepared MI-65 for access to investigate dehumidifier problems. During the evening shift, Operations resumed sending beam to NuMI at 5:41 PM; the MI-65 dehumidifier was left off. Operations disabled beam to T-1015 at 6:02 PM. At 10:14 PM, Operations reported a power glitch that tripped off a number of devices, but the stack, stash, and store remained untouched. Everything reset. Saturday July 2 The midnight shift began with Operations monitoring store 8855 (79.95E30) and the stash (269.30E10), with Pbar stacking (15.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-1015 at 4:29 AM. Operations terminated store 8855 at 5:42 AM, put the TEV into shot setup, and established store 8856 at 7:08 AM with an AIL of 366.79E30. During the day shift, at 1:28 PM, operators reported a high temperature for the MI-12 service building. FESS personnel investigated and reported that the portable air conditioner’s (AC) coil had froze, the other AC unit had tripped off, and the exhaust fan was off. MiniBooNE personnel had their beam turned off at 3:15 PM due to the high temperature. FESS personnel worked on the AC units and fan. During the evening shift, Operations resumed sending beam to NuMI at 5:41 PM. Operations disabled beam to T-1015 at 6:22 PM. Operations resumed sending beam to MiniBooNE at 8:19 PM. Operators reported that the P1 transfer line beam permit tripped off at 10:53 PM due to a power supply’s output (I:HT702) drifting low; everything reset. Operators reported short downtimes due to a Linac RF station (LRF2), a Linac Klystron RF station (KRF1), and a Booster RF station (BRF7). Sunday July 3 The midnight shift began with Operations monitoring store 8856 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8856 at 12:39 AM. Operators reported that a vacuum burst in the A1 transfer line propagated into the P1 transfer line and into the TEV. Operations put the TEV into shot setup and established store 8857 at 3:46 AM with an AIL of 357.41E30. Operations established beam to T-1015 at 4 AM. During the day shift, at 3:31 PM, a MI kicker (I:KPS5S) tripped off and wouldn’t reset until operators made an adjustment to its reservoir voltage. During the evening shift, at 4:56 PM, LRF2 tripped off many times. Operators contacted a Linac person. Operations put Pbar into standby mode and disabled beam to T-1015. Per Linac person instructions, operators made adjustments to LRF2 and Linac resumed sending beam at 6:01 PM. Operations terminated store 8857 at 9:11 PM, put the TEV into shot setup, and established store 8858 at 10:15 PM with an AIL of 363.58E30. At 11:31 PM, operators reported at that a MI Lambertson tripped off on a phase imbalance. Operators called in Lambertson personnel. Operators reported short downtimes due to KRF1, BRF7, BRF9, and BRF10 vacuum problems, and to KPS5S, E:QF107, BRF8, and MIRF15. Monday July 4 The midnight shift began with Operations monitoring store 8858, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, operators reported a MI-30 vacuum burst at 12:04 AM; everything reset. The MI Lambertson personnel reported, at 1:02 AM, finding and replacing a bad fan and a failed phase relay, which he bypassed. Operations established beam to T-1015 at 4:15 AM. A Linac quadrupole power supply’s breaker tripped off at 5:03 AM; it reset. Operators reported minor downtimes due to Booster vacuum pumps. The Plans for Monday The plans for today are to are to stack, stash, and store. The Run Coordinator will allow cryo system personnel to tighten the belts on the TEV sector A1 wet engine. |
|
More Information Comments and Suggestions |
| last modified 07/05/2011 email Fermilab |
| Security, Privacy, Legal | |